OpenStudy (anonymous):

how do i solve 4-(-6)/-3

OpenStudy (anonymous):

can you tell me how you got that?

OpenStudy (anonymous):

the two negatives become positives so it becomes 10 so 10/3 = 3 1/3

OpenStudy (anonymous):

my answer key says that the answer is -2

OpenStudy (anonymous):

Oh, I know what's happening. Use PEMDAS. The division has to happen first, so it's (-6)/(-3), which equals 2. 4-2 is 2.

OpenStudy (anonymous):

sorr maybe i should repost this... but the the problem is actually 4-(-6)/-5

OpenStudy (anonymous):

It'd be the same thing. Divide -6 by -5, and then subtract that answer from 4.

OpenStudy (anonymous):

-6/-5 = 1 1/5 so 4 - 1 1/5 = 2 4/5 = 2.8

OpenStudy (anonymous):

-6/-5 = 1 1/5? im sorry but can you explain how you go that answer this is new to me...

OpenStudy (anonymous):

when you divide two negatives you get a positive so it would actually be 6/5 then 5 goes into 6 once with 1/5 left over which gives you 1 1/5

OpenStudy (anonymous):

ohhh right improper fractions!!!

OpenStudy (anonymous):

yup lol

OpenStudy (anonymous):

what about 4- 1 1/5 how do you subtract those?

OpenStudy (anonymous):

If you change it to improper fractions it's easier, so instead of 4 it would be 20/5, and instead of 1&1/5 it would be 6/5. Then just subtract 20-6

OpenStudy (anonymous):

^ what she said lol

OpenStudy (anonymous):

the 20-6 = 14 14/5 = 2 4/5

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

only thing is that this asnwer key is saying that the answer is -2????

OpenStudy (anonymous):

then they aren't doing PEMDAS. so then 4- (-6) would equal 10 because the negatives become positive making it 4 + 6 10/ -5 would be -2 because the one negative makes the whole fraction negative

OpenStudy (anonymous):

omg so much easier!!!! Thanks nicole you are my new fan!!! :)

OpenStudy (anonymous):

i dont even know what PEMDAS is!!!

OpenStudy (anonymous):

Sarah i also appreciate your response!!! :)

OpenStudy (anonymous):

you'll learn pemdas in the future, but no problemo!

OpenStudy (anonymous):

I have another question...

OpenStudy (anonymous):

how would you get the square root of something w/out using a calculator????

OpenStudy (anonymous):

what's the problem?

OpenStudy (anonymous):

actually how do you even get a square root of something?

OpenStudy (anonymous):

example.... \[\sqrt{2} X \sqrt{15}=\]

OpenStudy (anonymous):

a square root is a number that when multiplied by itself equals a given number so like the square root of 9 would be 3 because 3*3 = 9 and the square root of 16 would be 4 because 4*4 = 16

OpenStudy (anonymous):

but 2 an\[\sqrt{30}\]d 5 dont have square roots.. could the anser to the above be

OpenStudy (anonymous):

woa! I messed that up...what i meant was that if 2 and 5 dont have square roots could the anser be \[\sqrt{30}\]

OpenStudy (anonymous):

**15 not 5